OpenStudy (anonymous):

Identify the slope of the asymptotes of 5(y -1)2 - 20(x - 1)2 =64

OpenStudy (jmark):

Given 5(y -1)2 - 20(x - 1)2 =64 => 5(y -1)^2/64 - 20(x – 1)^2/64 =1 =>5(y -1)^2/8^2-20(x – 1)^2/8^2 =1 =>(y -1)^2/(8/√5)^2 - (x – 1)^2/(8/√20)^2 =1 we have (y-1)^2/b^2-(x-1)^2/a^2=1 where slope of the asymptotes =±b/a =>slope of the asymptotes =± (8/√5)/(8/√20) => slope of the asymptotes =±√20/√5 Learn more about Find Asymptotes at http://goo.gl/axfC7w

OpenStudy (anonymous):

THANKS